OpenStudy (anonymous):

Hanif is Lusca’s son. Lusca is 5 times as old- as Hanif. By the time Hanif is 18, Lusca will be 8 times as old as Hanif is now. How old is Lusca?

OpenStudy (anonymous):

8 times or 8 years older?

OpenStudy (dumbcow):

L = 5H 8H = L+(18-H) 8H = 4H +18 H = 4.5 --> L = 22.5